OpenStudy (anonymous):

If you save one penny on January 1, two pennies on January 2, three pennies on January 3, and continue this pattern for one year (not a leap year), what will be the value of your entire savings, in dollars, at the end of that one year? Express your answer as a decimal. 667.95 543.09 768.97 537.09

OpenStudy (anonymous):

i said before \[\sum_{1}^i = [n*(n+1)]/2\] so use this fore the month s and solve it n is number of the days ex 30 0r 29 0r...

OpenStudy (anonymous):

667.95

OpenStudy (ranga):

Correct!

OpenStudy (anonymous):

did u solved?

OpenStudy (anonymous):

can you help me with my last question im having trouble on?

OpenStudy (anonymous):

Assume f(x) = 2x + 8 and g(x) = 5. What is the value of (f + g)(3)?

OpenStudy (anonymous):

ok it means f(x) + g(x) so f(3)= 2(3) +8 and g(x) always have a constant value so f(3)=14 g(3)=5 14+5=19

OpenStudy (anonymous):

thanks your awesome

OpenStudy (anonymous):

you r welcome